OpenStudy (anonymous):

(03.07) Which of the following is a solution of x2 − 7x = −5? 7 plus or minus the square root of 69 divided by two 7 plus or minus the square root of 29 divided by two. negative 7 plus or minus the square root of 29 divided by two negative 7 plus or minus the square root of 69 divided by two.

OpenStudy (welshfella):

x^2 -7x + 5 = 0 the general form is ax^2 + bx + c = 0 to solve for x use the quadratic formula x = [-b +/- (b^2 - 4ac)]/2a

OpenStudy (welshfella):

plug in the values a = 1, b = -7 and c = 5

OpenStudy (anonymous):

would it be negative 7 plus or minus the square root of 29 divided by two?

OpenStudy (welshfella):

No the first part is incorrect substituting for b = -7 -b = - (-7) = ?

OpenStudy (anonymous):

7

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

so B is the correct option

OpenStudy (anonymous):

i knew it!